# An algorithm for estimating the crossing number of dense graphs, and continuous analogs of the crossing and rectilinear crossing numbers

### v2

@@ Abstract
We present a deterministic $n^{2+o(1)}$-time algorithm that approximates the crossing number of any graph $G$ of order $n$ up to an additive error of $o(n^{4})$. We also provide a randomized polynomial-time algorithm that constructs a drawing of $G$ with $\text{cr}(G)+o(n^{4})$ crossings. These results are made interesting by the well known fact that every dense $n$-vertex graph has crossing number $\Theta(n^{4})$. Our work builds on a technique developed by Fox, Pach and Súk [18], who obtained very similar results for the rectilinear crossing number.
The results in [18] and in this paper imply that the (normalized) crossing and rectilinear crossing numbers are estimable parameters. Motivated by this, we introduce two graphon parameters, the crossing density and the rectilinear crossing density, and then we prove that, in a precise sense, these are the correct continuous analogs of the crossing and rectilinear crossing numbers of graphs.
@@ 1 Introduction
We work only with finite, simple and undirected graphs.
Let $G=(V,E)$ be a graph. A drawing of $G$ is a representation in which the vertices are mapped to distinct points on the plane and the edges are represented by simple continuous curves. We further assume that no edge goes through a vertex other than its endpoints, no two edges are tangent at any point, and no three of them have an interior point in common. A crossing is a common interior point of two edges in a drawing. The crossing number of $G$, denoted by $\mathop{\mathrm{cr}}(G)$, is the minimum number of crossing points between edges when the minimum is taken over all drawings of $G$. Note that any drawing of $G$ with $\mathop{\mathrm{cr}}(G)$ crossings has the additional property that no two edges cross more than once and no two adjacent edges cross. A straight-line drawing of $G$ is a drawing such that each edge is represented by a segment joining the corresponding endpoints. The rectilinear crossing number of $G$, $\overline{\mathop{\mathrm{cr}}}(G)$, is the least number of crossings amongst all straight-line drawings of $G$. Clearly, $\mathop{\mathrm{cr}}(G)\leq\overline{\mathop{\mathrm{cr}}}(G)$, and it is known that there are graphs for which the inequality is strict (a rather surprising example was obtained by Bienstock and Dean [8], who constructed graphs with crossing number $4$ but arbitrarily large rectilinear crossing number).
@@
exists (cf. [38]), determining it has proven to be very challenging. Currently, the best known bounds place this quantity between $0.379972$ and $0.380473$, these are due to Ábrego et al. [1] and Fabila-Monroy and López [17], respectively. For the crossing number, it is conjectured that
@@
exists and, by the result we just mentioned, is bounded from above by $3/8$. An interesting consequence of this is that this limit differs from the one for rectilinear crossing numbers. The asymptotic behaviours of $\mathop{\mathrm{cr}}(K_{m,n})$ and $\overline{\mathop{\mathrm{cr}}}(K_{m,n})$ are not completely understood either (see [37] and the references therein). The rectilinear crossing number of $K_{n}$ is closely related to $k$-sets and $k$-edges [^2] Given a finite set $P$ of points on the plane, a $k$-set is a $k$-element subset $S\subset P$ for which there exists a half-plane $H$ with $H\cap P=S$, and a $k$-edge consists of a pair of points $p,q\in P$ such that one of the closed half-planes determined by the line trough $p$ and $q$ contains exactly $k$ points of $P$. [31, 4], as well as to Sylvester’s four point problem [^3] Sylvester’s four point problem asks for the probability that four points chosen at random from a region $R$ of the plane lie in convex position. When $R$ is the whole plane, there are a few natural distributions from which one can choose the points; this led to some different answers being published. See [36] for an overview of the history of the problem. [38]. We shall come back to the later of this connections in section 4.
@@
Moving on to the computational aspects of the problem, computing the crossing number is known to be NP-complete [20], while determining the rectilinear crossing numbers is complete for the existential theory of reals [7], and hence NP-hard. For any fixed $k$, there is a linear time algorithm that decides whether $\mathop{\mathrm{cr}}(G)\leq k$ (i.e., the crossing number is fixed parameter tractable). In contrast, it is NP-hard to determine if $\overline{\mathop{\mathrm{cr}}}(G)\leq k$ holds. A considerable amount of work has been put into developing approximation algorithms for both $\mathop{\mathrm{cr}}(G)$ and $\overline{\mathop{\mathrm{cr}}}(G)$. A graph drawing technique of Bhatt and Leighton [6], in conjunction with the the results of Leighton and Rao [27] on balanced cuts, can be combined to find, in polynomial time, a straight-line drawing of any bounded degree $n$-vertex graph $G$ with no more than $O(\log^{4}n(n+\mathop{\mathrm{cr}}(G))$ crossings. This was later improved to $O(\log^{3}n(n+\mathop{\mathrm{cr}}(G))$ by Even et al. [16], and then to $O(\log^{2}n(n+\mathop{\mathrm{cr}}(G))$ as a result of the improved approximation algorithm for optimal balanced cuts by Arora et al. [3]. The major drawback of the aforementioned algorithm is that it provides no information in the case that $\mathop{\mathrm{cr}}(G)$ is at most linear in $n$. It wasn’t until several years later that Chuzhoy [12], using the edge planarization method [^4] A subset of the edges of a graph is called planarizing if deleting each of its elements results in a planar graph. The method can be summarized as follows: First, one tries to find a small planarizing set of edges and computes a planar drawing of the graph that we get after deleting those edges. Then, the edges of the planarizing set are carefully added to the drawing one by one until we obtain a drawing of the original graph. from [14], found a polynomial-time $O(n^{9/10})$-approximation algorithm for $\mathop{\mathrm{cr}}(G)$ for bounded degree graphs. Building further on the edge planarization method, Kawarabayashi and Sidiropoulos [24, 25] improved the approximation ratio to $O(n^{1/2})$, and then Mahabadi and Tan [13] found a randomized $O(n^{1/2-\delta})$-approximation algorithm, where $\delta>0$ is a constant.
@@
The celebrated crossing lemma, discovered simultaneously by Ajtai at al. [2] and Leighton [26], tells us that $\mathop{\mathrm{cr}}(G)\geq\frac{|E|^{3}}{64|V|^{2}}$, so long as $|E|\geq 4|V|$. An immediate consequence of this result is that if $G$ is dense (i.e., it has $\Omega(|V|^{2})$ edges) then both $\mathop{\mathrm{cr}}(G)$ and $\overline{\mathop{\mathrm{cr}}}(G)$ are $\Omega(|V|^{4})$. Fox, Pach and Suk [18] presented an algorithm that constructs a straight-line drawing of $G$ with $\overline{\mathop{\mathrm{cr}}}(G)+o(|V|^{4})$ crossings. More precisely, they showed the following.
@@
There is a deterministic $n^{2+o(1)}$-time algorithm that computes a straight-line drawing of any given graph $n$-vertex graph $G$ with no more than
@@
Note that if $G$ is dense then the number of crossings in the drawing provided by this algorithm is $(1+o(1))\overline{\mathop{\mathrm{cr}}}(G)$. In section 3, we obtain a similar result for the crossing number.
@@
There exists a deterministic $n^{2+o(1)}$-time algorithm that for any given graph $n$-vertex graph $G$ approximates $\mathop{\mathrm{cr}}(G)$ up to an additive error of $O(n^{4}/(\log\log n)^{\delta^{\prime}})$. Furthermore, there is a randomized polynomial-time algorithm that, with probability $1-o(1)$, computes a drawing of $G$ with
@@
At a high level, this algorithm follows the strategy devised by Fox, Pach and Suk for the proof of Theorem 1.1. The main novel ingredient in our proof is Theorem 3.3, which provides a bound on the difference between the crossing numbers of two graphs of the same order in terms of their cut distance.
A consequence of the results in [18] and in Section 3 of this paper is that the (normalized) crossing and rectilinear crossing numbers are estimable/testable parameters, in the sense of [9]. Motivated by this, in Section 4 we define two graphon parameters, which we call the crossing density and the rectilinear crossing density, by means of continuous analogs of the notions of drawings and straight-line drawings of graphs. We show that both of these parameters are continuous with respect to the cut norm and that, in a precise sense, they behave as the limits of the crossing number and the rectilinear crossing number of graphs. This fact is related to some of the problems we mentioned earlier about the asymptotic behaviours of crossing numbers. Some other properties of the crossing and rectilinear crossing densities of graphons are also discussed in this section. We hope that our results might prove useful in the study of crossing and rectilinear crossing numbers of dense graphs. The basic definitions regarding graphons, cut distance and estimable parameters are included in the following section.
@@ 2 Preliminaries
By an edge weighted graph, we mean a graph $G=(V,E)$ where each edge $(u,v)\in E$ has a weight $w_{G}(u,v)\in[0,1]$ assigned to it. We write $w_{G}(u,v)=0$ whenever $(u,v)\notin E$. We shall work mostly with edge weighted graph, and we often refer to them simply as graphs.
2.1 Crossing numbers of edge weighted graphs
Following [18], we extend the definitions of the crossing and the rectilinear crossing numbers to edge weighted graphs. For an edge weighted graph $G(V,E)$ and a drawing $\mathcal{D}$ of $G$, let $C(\mathcal{D})$ denote the multi-set of all pairs of edges that cross each other in the drawing, with the proper multiplicity (i.e., if two edges cross each other at $k$ points then this pair appears $k$ times in $C(\mathcal{D})$). Now, let
@@ 2.1 Crossing numbers of edge weighted graphs
and define $\mathop{\mathrm{cr}}(G)$ as the least value of $\mathop{\mathrm{cr}}(G,\mathcal{D})$ over all drawings of $G$. Similarly, $\overline{\mathop{\mathrm{cr}}}(G)$ is the minimum of $\mathop{\mathrm{cr}}(G,\mathcal{D})$ where $\mathcal{D}$ ranges over all straight-line drawings of $G$. We say that a drawing $\mathcal{D}$ of $G$ attains $\mathop{\mathrm{cr}}(G)$ if $\mathop{\mathrm{cr}}(G,\mathcal{D})=\mathop{\mathrm{cr}}(G)$, and that it attains $\overline{\mathop{\mathrm{cr}}}(G)$ if it is a straight-line drawing and $\mathop{\mathrm{cr}}(G,\mathcal{D})=\overline{\mathop{\mathrm{cr}}}(G)$. A drawing such that no two edges cross more than once and no two adjacent edges cross will be called simple. Although it is not as self evident as in the unweighted case, one can readily show that there must exist a simple drawing which attains $\mathop{\mathrm{cr}}(G)$. Note that, for unweighted graphs, we can assign a weight of $1$ to every edge in order to recover the definitions of crossing number and rectilinear crossing number provided in the introduction.
We will need the following simple result, which fulfills the same role as Lemma 2 in [18].
@@
Let $G$ be an edge weighted graph on $n$ vertices such that the weight of each edge can be represented using no more than $B$ bits. Then we can find a drawing of $G$ that attains $\mathop{\mathrm{cr}}(G)$ in $2^{O(n^{4}\log n)}+2^{O(n^{2})}B^{2}$ time.
@@
Given a set $C\subset\binom{E}{2}$, we can determine whether there is a simple drawing $\mathcal{D}$ of $G$ with $C(\mathcal{D})=C$ in the following way: For any given edge $e\in E$, there are less than $n^{2}!$ possible orders in which the points where $e$ crosses the other edges may be arranged along the curve representing $e$. For a fixed order of the crossings along every edge, we can test whether it arises from an actual drawing of $G$ by placing a dummy vertex at each crossing and then using a planarity testing algorithm. Thus, we can test whether $C$ actually comes from a drawing in $n^{2}!^{O(n^{2})}=2^{O(n^{4}\log n)}$ time.
If we are given $C(\mathcal{D})$, then we can compute $\mathop{\mathrm{cr}}(G,\mathcal{D})$ in $O(n^{2}B^{2})$ time. Hence, we can find a simple drawing $\mathcal{D}$ of $G$ that attains $\mathop{\mathrm{cr}}(G)$ in $2^{O(n^{4}\log n)}+O(2^{\binom{n}{2}}n^{2}B^{2})$ time. ∎
In order to study graphons in section 4, it will be convenient to have a normalized version of the crossing and rectilinear crossing numbers. With this in mind, for an edge weighted graph $G$ on $n$ vertices we define its crossing density as $\mathop{\mathrm{cd}}(G)=\mathop{\mathrm{cr}}(G)/n^{4}$ and its rectilinear crossing density as $\overline{\mathop{\mathrm{cd}}}(G)=\overline{\mathop{\mathrm{cr}}}(G)/n^{4}$.

### v4

@@ Abstract
We present a deterministic $n^{2+o(1)}$-time algorithm that approximates the crossing number of any graph $G$ of order $n$ up to an additive error of $o(n^{4})$. We also provide a randomized polynomial-time algorithm that constructs a drawing of $G$ with $\text{cr}(G)+o(n^{4})$ crossings. These results yield a $1+o(1)$ approximation algorithm for the crossing number of dense graphs. Our work complements a paper of Fox, Pach and Súk [21], who obtained similar results for the rectilinear crossing number.
The results in [21] and in this paper imply that the (normalized) crossing and rectilinear crossing numbers are estimable parameters. Motivated by this, we introduce two graphon parameters, the crossing density and the rectilinear crossing density, and we prove that, in a precise sense, these are the correct continuous analogs of the crossing and rectilinear crossing numbers of graphs.
@@ 1 Introduction
We work with finite, simple and undirected graphs.
Let $G=(V,E)$ be a graph. A drawing of $G$ is a representation in which the vertices are mapped to distinct points on the plane and the edges are represented by simple continuous curves connecting their respective endpoints. We further assume that no edge goes through a vertex other than its endpoints, no two edges are tangent at any point, and no three of them have an interior point in common. A crossing is a common interior point of two edges in a drawing. The crossing number of $G$, denoted by $\mathop{\mathrm{cr}}(G)$, is the minimum number of crossing points between edges when the minimum is taken over all drawings of $G$. Note that any drawing of $G$ with $\mathop{\mathrm{cr}}(G)$ crossings has the additional property that no two edges cross more than once and no two adjacent edges cross. A straight-line drawing of $G$ is a drawing such that each edge is represented by a segment joining the corresponding endpoints. The rectilinear crossing number of $G$, $\overline{\mathop{\mathrm{cr}}}(G)$, is the least number of crossings amongst all straight-line drawings of $G$. Clearly, $\mathop{\mathrm{cr}}(G)\leq\overline{\mathop{\mathrm{cr}}}(G)$, and it is known that there are graphs for which the inequality is strict (a rather surprising example was obtained by Bienstock and Dean [8], who constructed graphs with crossing number $4$ but arbitrarily large rectilinear crossing number).
@@
exists (cf. [46]), finding it has proven to be very challenging. Currently, the best known bounds place this quantity between $0.379972$ and $0.380473$; these are due to Ábrego et al. [1] and Fabila-Monroy and López [19], respectively. For the crossing number, it is conjectured that
@@
exists and, as evidenced by the aforementioned constructions, is bounded from above by $3/8$. An interesting consequence of this is that this limit differs from the one for rectilinear crossing numbers. The asymptotic behaviours of $\mathop{\mathrm{cr}}(K_{m,n})$ and $\overline{\mathop{\mathrm{cr}}}(K_{m,n})$ are not completely understood either (see [45] and the references therein). The rectilinear crossing number of $K_{n}$ is closely related to $k$-sets and $k$-edges [^2] Given a finite set $P$ of points on the plane, a $k$-set is a $k$-element subset $S\subset P$ for which there exists a half-plane $H$ with $H\cap P=S$, and a $k$-edge consists of a pair of points $p,q\in P$ such that one of the closed half-planes determined by the line trough $p$ and $q$ contains exactly $k$ points of $P$. [36, 4], as well as to Sylvester’s four point problem [^3] Sylvester’s four point problem asks for the probability that four points chosen at random from a region $R$ of the plane lie in convex position. When $R$ is the whole plane, there are a few natural probability distributions from which one can choose the points; this led to some different answers being published. See [44] for an overview of the history of the problem. [46]. We shall come back to the later of these connections at the end of Section 4.
@@
Moving on to the computational aspects of the problem, computing the crossing number is known to be NP-complete [23], while determining the rectilinear crossing number is complete for the existential theory of reals [7] (and hence NP-hard). In fact, there is some $c>0$ such that approximating the crossing number of $G$ up to a factor of $1+c$ is NP-hard, even for cubic graphs [10]. However, for any fixed $k$, there is a linear time algorithm that decides whether $\mathop{\mathrm{cr}}(G)\leq k$ [27] (in particular, the crossing number is fixed parameter tractable). A considerable amount of work has been put into developing approximation algorithms for both $\mathop{\mathrm{cr}}(G)$ and $\overline{\mathop{\mathrm{cr}}}(G)$. A graph drawing technique of Bhatt and Leighton [6], in conjunction with the results of Leighton and Rao [32] on balanced cuts, can be combined to find, in polynomial time, a straight-line drawing of any bounded degree $n$-vertex graph $G$ with no more than $O(\log^{4}n(n+\mathop{\mathrm{cr}}(G))$ crossings. This was later improved to $O(\log^{3}n(n+\mathop{\mathrm{cr}}(G))$ by Even et al. [18], and then to $O(\log^{2}n(n+\mathop{\mathrm{cr}}(G))$ as a result of the improved approximation algorithm for optimal balanced cuts by Arora et al. [3]. It wasn’t until several years later that Chuzhoy [13], using the edge planarization method [^4] A subset of the edges of a graph is called planarizing if deleting each of its elements results in a planar graph. The method can be summarized as follows: First, one tries to find a small planarizing set of edges and computes a planar drawing of the graph that we get after deleting those edges. Then, the edges of the planarizing set are carefully added to the drawing one by one until we obtain a drawing of the original graph. from [15], found a polynomial-time $O(n^{9/10})$-approximation algorithm for $\mathop{\mathrm{cr}}(G)$ for bounded degree graphs. Building further on the edge planarization method, Kawarabayashi and Sidiropoulos [28, 29] improved the approximation ratio to $O(n^{1/2})$, and then Mahabadi and Tan [14] found a randomized $O(n^{1/2-\delta})$-approximation algorithm, where $\delta>0$ is a constant.
@@
The celebrated crossing lemma, discovered simultaneously by Ajtai at al. [2] and Leighton [31], tells us that $\mathop{\mathrm{cr}}(G)\geq\frac{|E|^{3}}{64|V|^{2}}$, so long as $|E|\geq 4|V|$. An immediate consequence of this result is that if $G$ is dense (that is, it has $\Omega(|V|^{2})$ edges) then both $\mathop{\mathrm{cr}}(G)$ and $\overline{\mathop{\mathrm{cr}}}(G)$ are $\Omega(|V|^{4})$. Fox, Pach and Suk [21] presented an algorithm that constructs a straight-line drawing of $G$ with $\overline{\mathop{\mathrm{cr}}}(G)+o(|V|^{4})$ crossings. More precisely, they showed the following.
@@
There is a deterministic $n^{2+o(1)}$-time algorithm that computes a straight-line drawing of any given $n$-vertex graph $G$ with no more than
@@
Note that if $G$ is dense then the number of crossings in the drawing provided by this algorithm is $(1+o(1))\overline{\mathop{\mathrm{cr}}}(G)$. We obtain a similar result for the crossing number.
@@
There exists a deterministic $n^{2+o(1)}$-time algorithm that for any given $n$-vertex graph $G$ approximates $\mathop{\mathrm{cr}}(G)$ up to an additive error of $O(n^{4}/(\log\log n)^{\delta^{\prime}})$. Furthermore, there is a randomized polynomial-time algorithm that, with probability $1-o(1)$, computes a drawing of $G$ with
@@
At a high level, the approximation part of the algorithm follows the strategy devised by Fox, Pach and Suk for the proof of Theorem 1.1. The main novel ingredient that is required for the proof of correctness of this algorithm is Theorem 3.3, which provides a bound on the difference between the crossing numbers of two graphs of the same order in terms of their distance in the cut norm.
A consequence of the results in [21] and Theorem 3.3 is that the (normalized) crossing and rectilinear crossing numbers are estimable/testable parameters, in the sense of [9]. Motivated by this, we define two graphon parameters, which we call the crossing density and the rectilinear crossing density, by means of continuous analogs of the notions of drawings and straight-line drawings of graphs. We show that both of these parameters are continuous with respect to the cut norm and that, in a precise sense, they behave as the limits of the crossing number and the rectilinear crossing number of graphs. This discussion is directly tied to some of the problems we mentioned earlier about the asymptotic behaviors of crossing numbers. We hope that our results might prove useful in the study of crossing and rectilinear crossing numbers of dense graphs.
Outline of the paper
The basic definitions regarding graphons, cut distance and estimable parameters, as well as some other necessary preliminaries, are included in Section 2. In Section 3, we prove Theorem 1.2; most of this section is in fact devoted to the proof of Theorem 3.3, which was mentioned above. The crossing and rectilinear crossing densities of graphons are defined and studied in Section 4. Lastly, we discuss some unanswered questions in Section 5.
@@ 2 Preliminaries
By an edge-weighted graph, we mean a graph $G=(V,E)$ where each edge $(u,v)\in E$ has a weight $w_{G}(u,v)\in[0,1]$ assigned to it. We write $w_{G}(u,v)=0$ whenever $(u,v)\notin E$. We shall work mostly with edge-weighted graph, and we often refer to them simply as graphs.
2.1 Crossing numbers of edge-weighted graphs
Next, we extend the definitions of the crossing and the rectilinear crossing numbers to edge-weighted graphs. For an edge-weighted graph $G(V,E)$ and a drawing $\mathcal{D}$ of $G$, let $C(\mathcal{D})$ denote the multi-set of all pairs of edges that cross each other in the drawing, with the proper multiplicity (i.e., if two edges cross each other at $k$ points then this pair appears $k$ times in $C(\mathcal{D})$). Now, let
@@ 2.1 Crossing numbers of edge-weighted graphs
and define $\mathop{\mathrm{cr}}(G)$ as the least value of $\mathop{\mathrm{cr}}(G,\mathcal{D})$ over all drawings of $G$. Similarly, $\overline{\mathop{\mathrm{cr}}}(G)$ is the minimum of $\mathop{\mathrm{cr}}(G,\mathcal{D})$ where $\mathcal{D}$ ranges over all straight-line drawings of $G$. We say that a drawing $\mathcal{D}$ of $G$ attains $\mathop{\mathrm{cr}}(G)$ if $\mathop{\mathrm{cr}}(G,\mathcal{D})=\mathop{\mathrm{cr}}(G)$, and that it attains $\overline{\mathop{\mathrm{cr}}}(G)$ if it is a straight-line drawing and $\mathop{\mathrm{cr}}(G,\mathcal{D})=\overline{\mathop{\mathrm{cr}}}(G)$. A drawing such that no two edges cross more than once and no two adjacent edges cross will be called simple. Although it is not as self-evident as in the unweighted case, one can readily show that there must exist a simple drawing which attains $\mathop{\mathrm{cr}}(G)$. Note that, for unweighted graphs, we can assign a weight of $1$ to every edge in order to recover the definitions of crossing number and rectilinear crossing number provided in the introduction.
We will need the following simple result, which fulfills the same role as Lemma 2 in [21]. It seems very likely that a result of this kind has explicitly appeared somewhere else already, but we have been unable to find it.
@@
Let $G$ be an edge-weighted graph on $n$ vertices such that the weight of each edge can be represented using no more than $B$ bits. Then we can find a drawing of $G$ that attains $\mathop{\mathrm{cr}}(G)$ in $2^{O(n^{4}\log n)}+2^{O(n^{2})}B^{2}$-time.
@@
Given a set $C\subset\binom{E}{2}$, we can determine whether there is a simple drawing $\mathcal{D}$ of $G$ with $C(\mathcal{D})=C$ in the following way: For any given edge $e\in E$, there are less than $n^{2}!$ possible orders in which the points where $e$ crosses the other edges may appear along the curve representing $e$. For a fixed order of the crossings along every edge, we can test whether it arises from an actual drawing of $G$ by placing a dummy vertex at each crossing and then using a linear time planarity testing algorithm (see [43]). Thus, we can test whether $C$ actually comes from a drawing in $n^{2}!^{O(n^{2})}=2^{O(n^{4}\log n)}$-time.
If we are given $C(\mathcal{D})$, then we can compute $\mathop{\mathrm{cr}}(G,\mathcal{D})$ in $O(n^{2}B^{2})$-time. Note also that the number of possible $C$’s we must check is $2^{|E|}\leq 2^{\binom{n}{2}}$. Hence, we can find a simple drawing $\mathcal{D}$ of $G$ that attains $\mathop{\mathrm{cr}}(G)$ in $2^{O(n^{4}\log n)}+O(2^{\binom{n}{2}}n^{2}B^{2})$-time. ∎
In order to study graphons in section 4, it will be convenient to have a normalized version of the crossing and rectilinear crossing numbers. With this in mind, for an edge-weighted graph $G$ on $n$ vertices we define its crossing density as $\mathop{\mathrm{cd}}(G)=\mathop{\mathrm{cr}}(G)/n^{4}$ and its rectilinear crossing density as $\overline{\mathop{\mathrm{cd}}}(G)=\overline{\mathop{\mathrm{cr}}}(G)/n^{4}$.